OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5?

OpenStudy (anonymous):

Take the first equation, y = -2x + 5 We know that this has a negative gradient as the x-coefficient is negative. This means it will be travelling from top left to bottom right. We only need two points to plot a line, so we can use the x and y intercepts as it makes the maths easy substituting in 0! You could however, put in any number for x and obtain the y value to find a point on the line. Anyway, back to intercepts. When x = 0, y = -2(0) + 5 = 5 so the point (0, 5) is on the line. When y = 0, 0 = -2x + 5 2x = 5 x = 2/5 = 0.4 So the point (0.4, 0) is also on the line. Plot these 2 points and draw a line through them. Make the line as long as possible on your page. Now consider the second equation. We need to find two points again. When x = 0, y = (1/2)(0) - 5 y = -5 So, (0, -5) is a point on the line. When y = 0, 0 = (1/2)x - 5 (1/2)x = 5 x = 10. So (10, 0) is a point on the line. Plot these two points and draw a nice long line through them. Now, the lines should cross somewhere, and this is the point of intersection. Simply read it off and there's your answer.

OpenStudy (anonymous):

This is history not math. :P

OpenStudy (anonymous):

lol, oops :)

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

i got a medal anyway though :).